# Binary Hypothesis Testing-Based Low-Complexity Beamspace Channel Estimation for mmWave Massive MIMO Systems

### v1

@@ IV-A Simulation Configuration
To assess the performance of the proposed algorithm, we consider a BS with 256256256-element ULA employing all-digital beamforming, and 16 users equipped with single antennas in the cell. We assume the antenna spacing of BS is half-wavelength, and the antenna of each user is omnidirectional. For the channel models, we consider both synthetic data following the distribution in (9) and realistic channel models. The realistic channel models are generated for both LoS and non-LoS (NLoS) scenarios based on the QuaDRiGa mmMAGIC UMi[47] with a carrier frequency of 50 GHz. The generated channel models contain many near-zero elements but rarely include exact zeros, because the signals in realistic channel models arrive from arbitrary directions and do not align well with DFT bins, resulting in power leakage[48]. Therefore, directly defining the activity rate using the $\ell_{0}$-norm is infeasible. To quantify the beamspace sparsity in a realistic setting where channel coefficients are rarely exactly zero, we define the activity rate based on an energy thresholding criterion. Let $|\bar{\mathbf{h}}^{\textnormal{sort}}|^{2}=[|\bar{h}^{\textnormal{sort}}_{1}|^{2},\cdots,|\bar{h}^{\textnormal{sort}}_{M}|^{2}]^{T}\in\mathbb{R}^{M}$ denote the beamspace element-wise channel power vector which is sorted in descending order, i.e., $|\bar{h}^{\textnormal{sort}}_{1}|^{2}\geq\cdots\geq|\bar{h}^{\textnormal{sort}}_{M}|^{2}$. Then, for a given energy threshold $\eta$, the activity rate is defined as $q_{\textnormal{energy}}=A_{\eta}/M$, where the number of active beam $A_{\eta}$ is
@@
This means that the beam components that, even when combined among themselves, contribute less than $(1-\eta)$ of the total channel power are classified as inactive. For example, if $\eta=0.9$, the beams that constitute 90% of the total power in the noiseless beamspace channel are considered active, while the remaining smallest beams that represent the residual 10% are regarded as inactive. In this paper, we consider $\eta=0.99$, to neglect the 1% power leakage. [^1] The energy threshold $\eta$ can also be set to other values, e.g., 0.95, 0.999, or other appropriate values to regard near-zero elements as inactive. This reflects the effective sparsity of the channel by capturing how concentrated the energy is among a few dominant beams. This classification metric is also used to compute $P_{D}$ and $P_{FA}$. The cost parameter $C$ used in the hypothesis testing is set to 5. A detailed empirical justification for this choice is presented in a subsequent subsection.
For the simulations, we conducted 10000 Monte Carlo trials. To conduct performance evaluation of the proposed channel estimation algorithm, we compare its estimation accuracy with LS estimation and other baseline channel estimation algorithms. For baselines, we consider representative mmWave beamspace channel estimation algorithms that exploit sparsity: BEACHES[42] as a low-complexity denoiser, LAMP-GM[39] with the optimal number of Gaussian mixtures as a supervised learning-based method, and LDGEC-SURE[33] as an unsupervised learning-based estimator. Also, as a reference, the results for “Perfect detection” use the same noise-zeroing-based approach for denoising, but it is based on perfect detection of nonzero beamspace channel vector element. Note that this approach is equivalent to the theoretically ideal case of the proposed algorithm which ignores the estimation errors of activity rate, noise power, and SNR.
@@ IV-C Performance Evaluation
Fig. 1 presents the performance of activity rate estimation compared to the ground-truth value for synthetic and realistic channel data. For the synthetic data, the estimator achieves high accuracy when $\mathrm{SNR}\geq 5$ dB, while it tends to slightly overestimate $q$ under low-SNR conditions such as $\mathrm{SNR}=-5$ dB. As discussed in Theorem 2, the estimator shows relatively higher accuracy for smaller ground-truth $q$. Meanwhile, for realistic channel data, the estimation accuracy degrades compared to the synthetic case, as the realistic beamspace channels do not strictly follow the Bernoulli-complex Gaussian distribution. Under LoS conditions, the estimator tends to underestimate the activity rate, since the strong LoS component increases the fourth-order moment of the beamspace channel vector. Compared to NLoS, the LoS case exhibits smaller estimation variance but lower accuracy. In contrast, the NLoS case yields higher accuracy with a larger variance.
@@
[…]
Fig. 5 illustrates the MSE depending on the noise power estimation error. Note that noise power estimation error also propagate to the SNR estimate in (17), and the activity rate estimate in (24). In both cases, underestimating the noise power—i.e., when the noise power error is negative—does not critically affect the MSE. Slight overestimation also has negligible impact on the estimation performance across all SNR regimes. Even when the noise power is overestimated by nearly a factor of two, the MSE remains stable at 5 dB and 15 dB SNR levels. In contrast, at low SNR like -5 dB, such severe overestimation leads to a noticeable increase in MSE. Nevertheless, such cases rarely occur in practice, as such severe noise power overestimation is uncommon in low-SNR conditions[44].
@@
Fig. 6 illustrates the receiver operating characteristic (ROC) curve of signal element detection. As the SNR increases, the ROC curve shifts toward the upper-left corner, indicating a higher probability of detection at a given false alarm probability, which means the significant improvement of detection performance with increasing SNR. This trend is consistent among the theoretical predictions and measured results, and also in both LoS and NLoS scenarios. This trend is intutive: higher noise power makes detection more difficult, while lower noise power implies ease of detection due to larger differences between the magnitudes of noise-relevant and signal-dominant elements. In LoS scenarios, the measured results exhibit lower estimation performance compared to the theoretical ROC curve. On the other hand, measured result shows similar performance to the theoretical curve in NLoS scenarios. It is because LoS beamspace channel has a LoS path with dominant gain which has massive magnitude compared to other paths, which makes the approximation to Bernoulli-complex Gaussian prior less efficient. Comparing the curves, the ROC curve for the LoS case lies further toward the upper-left corner than that of the NLoS case, which reflects better detection capability. It is because LoS channel has more sparsity than NLoS, and stronger sparsity, i.e., smaller $q$, leads to improved detection probability, which is described in (27).
V Conclusion
@@ Appendix A Derivation of Hypothesis Testing Threshold
×$\quad\times\ \ln\left(\frac{\|\bar{\mathbf{h}}\|^{2}/p(\mathcal{H}_{1})+E_{0}}{E_{0}}\times\frac{1-p(\mathcal{H}_{1})}{p(\mathcal{H}_{1})}C\right).$
@@
×$\quad\times\ \ln\left(\left(\frac{\|\bar{\mathbf{h}}\|^{2}}{E_{0}p(\mathcal{H}_{1})}+\frac{E_{0}}{E_{0}}\right)\times\frac{1-p(\mathcal{H}_{1})}{p(\mathcal{H}_{1})}C\right)$
@@
×$\quad\times\ \ln\left(\left(\frac{1}{E_{0}}\frac{\|\bar{\mathbf{h}}\|^{2}}{p(\mathcal{H}_{1})}+1\right)\frac{1-p(\mathcal{H}_{1})}{p(\mathcal{H}_{1})}C\right)$
@@ Appendix B Derivation of Theoretical $P_{FA}$ and $P_{D}$
[…]
@@
[…]
@@ C-A Proof of Theorem 1
$\bar{h}_{m}^{\star}=\begin{cases}0&\textnormal{with probability }1-P_{FA},\\ \bar{e}_{m}&\textnormal{with probability }P_{FA}.\\ \end{cases}$ (53)
@@ C-D Proof of Theorem 2
$\textnormal{Var}(\hat{q})=\textnormal{Var}\left(\hat{q}_{u}+\tilde{q}\right)=\textnormal{Var}(\hat{q}_{u})+\textnormal{Var}(\tilde{q})+2\mathrm{Cov}(\hat{q}_{u},\tilde{q}).$ (63)
@@
[…]
@@
Meanwhile, utilizing the inequality, the variation of $|\tilde{q}|$ is bounded as
@@
$\textnormal{Var}(\hat{q})$ $=\textnormal{Var}(\hat{q}_{u})+\textnormal{Var}(\tilde{q})+2\mathrm{Cov}(\hat{q}_{u},\tilde{q})$ (73)
@@
Also, this lower/upper bounds imply that $\textnormal{Var}(\hat{q})\propto q^{2}$. ∎

### v2

@@ IV-A Simulation Configuration
To assess the performance of the proposed algorithm, we consider a BS with $128$-element ULA employing all-digital beamforming, and 8 users equipped with single antennas in the cell. We assume the antenna spacing of BS is half-wavelength, and the antenna of each user is omnidirectional. For the channel models, we consider both synthetic data following the distribution in (8) and realistic channel models. The realistic channel models are generated for both LoS and non-LoS (NLoS) scenarios based on the QuaDRiGa mmMAGIC UMi[50] with a carrier frequency of 50 GHz. The generated channel models contain many near-zero elements but rarely include exact zeros, because the signals in realistic channel models arrive from arbitrary directions and do not align well with DFT bins, resulting in power leakage[51]. Therefore, directly defining the activity rate using the $\ell_{0}$-norm is infeasible. To quantify the beamspace sparsity in a realistic setting where channel coefficients are rarely exactly zero, we define the activity rate based on an energy thresholding criterion. Let $|\bar{\mathbf{h}}^{\textnormal{sort}}|^{2}=[|\bar{h}^{\textnormal{sort}}_{1}|^{2},\cdots,|\bar{h}^{\textnormal{sort}}_{M}|^{2}]^{T}\in\mathbb{R}^{M}$ denote the beamspace element-wise channel power vector which is sorted in descending order, i.e., $|\bar{h}^{\textnormal{sort}}_{1}|^{2}\geq\cdots\geq|\bar{h}^{\textnormal{sort}}_{M}|^{2}$. Then, for a given energy threshold $\eta$, the activity rate is defined as $q_{\textnormal{energy}}=A_{\eta}/M$, where the number of active beam $A_{\eta}$ is
@@
This means that the beam components that, even when combined among themselves, contribute less than $(1-\eta)$ of the total channel power are classified as inactive. For example, if $\eta=0.9$, the beams that constitute 90% of the total power in the noiseless beamspace channel are considered active, while the remaining smallest beams that represent the residual 10% are regarded as inactive. In this paper, we consider $\eta=0.99$, to neglect the 1% power leakage. [^1] The energy threshold $\eta$ can also be set to other values, e.g., 0.95, 0.999, or other appropriate values to regard near-zero elements as inactive. This reflects the effective sparsity of the channel by capturing how concentrated the energy is among a few dominant beams. This classification metric is also used to compute $P_{D}$ and $P_{FA}$. The cost parameter $C$ used in the hypothesis testing is set to 5, which shows a robust estimation performance. A detailed empirical justification for this choice is presented in a subsequent subsection.
For the simulations, we conducted 10000 Monte Carlo trials. To conduct performance evaluation of the proposed channel estimation algorithm, we compare its estimation accuracy baseline channel estimation algorithms. For the baseline, we consider least-squares (LS) estimation, which is known as a representative and basic pilot-based channel estimation method, and mmWave beamspace channel estimation algorithms that exploit sparsity: BEACHES[45] as a low-complexity denoiser, LAMP-GM[41] with the optimal number of Gaussian mixtures as a supervised learning-based method, and LDGEC-SURE[35] as an unsupervised learning-based estimator. Also, as a reference, the results for “Perfect detection” use the same noise-zeroing-based approach for denoising, but it is based on perfect detection of nonzero beamspace channel vector element. Note that this approach is equivalent to the theoretically ideal case of the proposed algorithm which ignores the estimation errors of activity rate, noise power, and SNR.
@@ IV-C Performance Evaluation
Fig. 1 presents the performance of activity rate estimation compared to the ground-truth value for synthetic and realistic channel data. For the synthetic data, the estimator achieves high accuracy when $\mathrm{SNR}\geq 5$ dB, while it tends to slightly overestimate $q$ under low-SNR conditions such as $\mathrm{SNR}=-5$ dB. As discussed in Theorem 2, the estimator shows relatively higher accuracy for a smaller ground-truth $q$. Meanwhile, for realistic channel data, the estimation accuracy degrades compared to the synthetic case, as the realistic beamspace channels do not strictly follow the Bernoulli-complex Gaussian distribution. The result shows lower estimation accuracy with larger variance in lower SNR conditions. Also, the LoS case exhibits a smaller estimation variance compared to that of the NLoS case.
@@
Fig. 4 illustrates the variation of MSE depending on the cost $C$. In both LoS and NLoS scenarios, the MSE tends to be higher when $C<1$, except in the NLoS case at the SNR of 15 dB, where the MSE remains nearly constant. This is because $C<1$ implies $C_{10}<C_{01}$, indicating that the testing is stricter against misdetection but more tolerant toward false alarms, which results in increased residual noise and degraded denoising performance for sparse signals. In LoS scenarios, increasing $C$ generally leads to a reduction in MSE or preserves it at a similar level, with the most notable improvements occurring when $C\leq 3$. In contrast, under NLoS conditions, increasing $C$ results in a noticeable MSE reduction only in the low-SNR case ($\text{SNR}=-5$ dB), while in higher SNR settings ($\text{SNR}\geq 5$ dB), a slight increase in MSE is observed. Nevertheless, this increase remains negligible. Overall, selecting $C$ around 5 does not significantly affect estimation performance across various conditions.
Fig. 5 illustrates the MSE depending on the noise power estimation error. Note that noise power estimation error also propagates to the SNR estimate in (16), and the activity rate estimate in (23). In both cases, underestimating the noise power—i.e., when the noise power error is negative—does not critically affect the MSE. Slight overestimation also has a negligible impact on the estimation performance across all SNR regimes. Even when the noise power is overestimated by nearly a factor of two, the MSE remains stable at 5 dB and 15 dB SNR levels. In contrast, at low SNRs (e.g., -5 dB), such severe overestimation results in a noticeable increase in the MSE. Nevertheless, such cases rarely occur in practice, as such severe noise power overestimation is uncommon in low-SNR conditions[47].
Fig. 6 illustrates the receiver operating characteristic (ROC) curve of signal element detection. The theoretical predictions and measured results show similar values, which implies that the Bernoulli-complex Gaussian model approximates the beamspace sparse channel well. As the SNR increases, the ROC curve shifts toward the upper-left corner, indicating a higher probability of detection at a given false alarm probability, which means the significant improvement of detection performance with increasing SNR. This trend is consistent among the theoretical predictions and measured results, and also in both LoS and NLoS scenarios. This trend is intutive: higher noise power makes detection more difficult, while lower noise power implies ease of detection due to larger differences between the magnitudes of noise-relevant and signal-dominant elements. In LoS scenarios, the measured results exhibit lower estimation performance compared to the theoretical ROC curve. On the other hand, measured result shows similar performance to the theoretical curve in NLoS scenarios. It is because LoS beamspace channel has a LoS path with dominant gain which has massive magnitude compared to other paths, which makes the approximation to Bernoulli-complex Gaussian prior less efficient. Comparing the curves, the ROC curve for the LoS case lies further toward the upper-left corner than that of the NLoS case, which reflects better detection capability. It is because LoS channel has more sparsity than NLoS, and stronger sparsity, i.e., smaller $q$, leads to improved detection probability, which is described in (26).
V Conclusion
@@ Appendix A Derivation of Hypothesis Testing Threshold
$\quad\times\ \ln\left(\frac{\|\bar{\mathbf{h}}\|^{2}/p(\mathcal{H}_{1})+E_{0}}{E_{0}}\times\frac{1-p(\mathcal{H}_{1})}{p(\mathcal{H}_{1})}C\right).$
@@
$\quad\times\ \ln\left(\left(\frac{\|\bar{\mathbf{h}}\|^{2}}{E_{0}p(\mathcal{H}_{1})}+\frac{E_{0}}{E_{0}}\right)\times\frac{1-p(\mathcal{H}_{1})}{p(\mathcal{H}_{1})}C\right)$
@@
$\quad\times\ \ln\left(\left(\frac{1}{E_{0}}\frac{\|\bar{\mathbf{h}}\|^{2}}{p(\mathcal{H}_{1})}+1\right)\frac{1-p(\mathcal{H}_{1})}{p(\mathcal{H}_{1})}C\right)$
@@ Appendix B Derivation of Theoretical $P_{FA}$ and $P_{D}$
$P_{FA}=\int_{\tau}^{\infty}\frac{1}{E_{0}}\exp\left(-\frac{t}{E_{0}}\right)dt=\exp\left(-\frac{\tau}{E_{0}}\right).$ (44)
@@
$P_{D}$ $=\int_{\tau}^{\infty}\frac{1}{E_{0}+\|\mathbf{h}\|^{2}/q}\exp\left(-\frac{t}{E_{0}+\|\mathbf{h}\|^{2}/q}\right)dt$ (47)
@@ C-A Proof of Theorem 1
[…]
@@ C-D Proof of Theorem 2
[…]
@@
$\sqrt{M}(\hat{\mu}_{4}-\mu_{4})\overset{D}{\longrightarrow}\mathcal{N}(0,\sigma^{2}),$ (65)
where $\sigma^{2}>0$ is an arbitrary constant. Thus, $\textnormal{Var}(\hat{\mu}_{4})=\sigma^{2}/M$. Also, $f(\mu)$ is differentiable in a neighborhood of $\mu_{4}$. Hence, delta method is applicable, and provides a valid first-order approximation of the variance of $\hat{q}_{u}$ as follows:
@@
[…]
@@
[…]
@@
Also, this lower/upper bounds imply that
$\textnormal{Var}(\hat{q})\propto q^{2}.$ (74)